OpenStudy (anonymous):

two cars leave city A at the sametime to travel to city B. One car travels at 72km/h and the other travels at 78 km/h. if the slower car arrives 2o mins after the faster car.. how far apart are city A and city B? pls help me answer dis

OpenStudy (anonymous):

u r in grade ????

OpenStudy (anonymous):

grade 10

OpenStudy (anonymous):

pls help me.

OpenStudy (anonymous):

Let A and B be s km apart Distance covered is velocity*time Car A is travelling at 72km/h So it takes car A 72x hours to reach city B Car B is travelling at 78km/h So it takes car B 78y hours to reach city B We know that x = y+1/3, because it takes car A 20 minutes longer (1/3 hours longer) And they cover the same distance so 72(y+1/3) = 78y 72y + 24 = 78y y = 4 And the distance is 78*4 = 312km

OpenStudy (nikvist):

\[t=\frac{s}{v_1}-\frac{s}{v_2}=s\cdot\frac{v_2-v_1}{v_1v_2}\] \[s=\frac{v_1v_2}{v_2-v_1}\cdot t=\frac{72\cdot 78}{78-72}\cdot\frac{1}{3}km=312\,km\]

OpenStudy (anonymous):

First find the distance between the two cars when the faster one has arrived at city B (which is): 72x20/60 = 24km difference Then, since the faster car drives 6km/hr faster than the slower one, to find the total time needed to get to city B with the faster car you: 24/6 = 4hrs Then you multiply this by the speed of the faster car: 4x78 = 312km